OpenStudy (anonymous):

Solve the equation -12 = 6(2x + 4)

OpenStudy (anonymous):

-12 = 6 ( 2 x + 4) -12 = 12 x + 24 -12-24 = 12x -36 = 12x x = -3

pooja195 (pooja195):

That's right.

OpenStudy (mrcoolguy):

x=-3